Give the Hardy–Weinberg expected genotypic frequencies for an autosomal locus with three alleles, where the frequencies of alleles A1, A2, and A3 are defined as p, q, and r, respectively. Set the problem up and show your work.

Tay–Sachs disease is an autosomal recessive disorder. Among Ashkenazi Jews, the frequency of Tay–Sachs disease is 1 in 3600. If the Ashkenazi population is mating randomly for the Tay–Sachs gene, what proportion of the population consists of heterozygous carriers of the Tay–Sachs allele?
Answer #1

Answer:

Tay-sachs frequency (tt) = q^2 = 1/3600 = 0.00028

Tay-sachs allele frequency(t) = q = SQRT of q^2 = 0.0167

Normal allele frequency (T)= p = 1- 0.0167 = 0.9833

Heterozygous normal frequency = Tt = 2* 0.9833 * 0.0167 = 0.03284